Felix drove 800 miles in x hours on Monday.
a)Write a rational expression for his average speed.

b)On Tuesday he drove for 6 hours at the same average speed. Write a rational expression for his distance on Tuesday.

Answer by Mathtut(3670)

speed is equal to the distance divided by time
:
a)avg speed=(800/x)mi/hr
:
b)distance = rate times time
:
d=(800*6)/x=4800/x